3a. Fill in the table below for a long position in S&P 500 futures contract with 250 multiplier. Initial Margin is S 35000 10 points S&P 500 Future Price 2555 Investment Value Gain (+) Loss (-) Balance as Margin Time 2 2535 2575
Answer #1

Please provide rating...

Its not mentioned in the question as to what price future is bought we are assuming that future is bought on Time 1 @ 2555
Time S&P 500 Future price Investment value Gain/(Loss) Balance as Margin
1 2555 638750 0 35000
2555*250
2 2535 633750 -5000 30000
2535*250 =633750-638750 =35000-5000
3 2575 643750 10000 40000
2575*250 =643750-633750 =30000+10000